A company has total assets of as December 31, 2018 of $37,000. The company has the following accounts (12/31/18 balances in parentheses): Cash (Unknown), Accounts Payable ($5,000). Accounts Receivable ($9,000), Supplies ($1,000), Equipment ($22,000), Retained Earnings (Unknown), and Common Stock ($11,000). The company has no other accounts. Solve for cash and Retained Earnings. Show your work.
Answer #1

total assets=cash+AR+Supplies+equipment

37000=Cash+9000+1000+22000

Cash=(37000-9000-1000-22000)=$5000

Total assets=total liabilities+Total equity

37000=AP+(Retained earnings+Common stock)

37000=5000+Retained earnings+11000

Hence Retained earnings=(37000-5000-11000)=$21000.